OpenStudy (anonymous):

Show work What NUMBER is 42% of 150 ?

OpenStudy (candy13106):

63

OpenStudy (mathstudent55):

Translate the question into math. "of" usually means "times" % = 1/100 = 0.01 An unknown number = x What number: x is: = 42% = 0.42 of: * 150: 150 x = 0.42 * 150 Just multiply 0.42 by 150